# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ 6.1.2 Numerical Results
We now present and analyze the results of each synthetic experiment. Figure 2 shows the normalized error metrics of each model on each function. The normalization is done within each column so that 1.00 and 0.00 represent, respectively, the highest and lowest error between all models for a specific metric and function, e.g., the MAE on McCormick’s function. The absolute values are also presented in Appendix B for additional information.
@@
The objective of Experiment A is to evaluate the ability of each model at learning over a highly adversarial training set while having the possibility to refine, indirectly through the validation phase, on a slightly out-of-sample distribution. This experiment was designed to see how different models’ regularization techniques may prevent them from overfitting on training outliers. As we can see, the WaDiRo-SCNN, regularized SCNN, and deep FNN perform similarly except when trying to learn Ackley where the WaDiRo-SCNN has the lowest test error. In this experiment, the linear models perform poorly compared to other models, we hypothesize that the validation phase is uncorrupted enough to let nonlinear models adjust to the non-convex functions during hyperparameter optimization. As expected the non-regularized SCNN leads to performance lower than its regularization counterpart in this particular setting.
@@ 6.2.1 Setting
In 2023, commercial and industrial sectors together represented 49.2% of Québec’s total energy consumption [51]. To accelerate the deployment of Québec’s virtual power plant (VPP), there is need for improving forecasting algorithms for individual non-residential buildings. These algorithms can be used for operations and planning decision-making but also for retroactively calculating rewards associated with shaved energy during demand response events [52]. Non-residential buildings comprise public buildings, e.g., public libraries, schools, and institutional buildings; commercial buildings, e.g., shops, cinemas, and offices; and industrial buildings, e.g., factories and warehouses.
Montréal’s continental humid climate delivers harsh winters and hot summers. As such, a significant amount of energy is consumed for both heating and cooling. Moreover, date and time features have a strong influence on non-residential buildings’ consumption patterns as their activities may change within days and weeks, e.g., opening hours, occupancy, and seasonal activities.
In this experiment, we seek to predict the year-round hourly energy consumption (kWh) of 16 non-residential buildings located in three adjacent distribution substations in Montréal, Québec, Canada. As we do not have any information on building activity, this task is done using only hourly consumption history and historical meteorological features. We use a 12-week history to train each model to predict the following week. We have approximately 2.5 years of collected data on each building. We proceed with a 12-week sliding window and collect the predictions for each new week. We use the data of the first year and a half for hyperparameter tuning and the data of the last year for performance assessment. The features and label are listed in Table 2.
@@
1 Day of the week, cosinus encoding $[0,1]$
2 Day of the week, sinus encoding $[0,1]$
3 Hour of the day, cosinus encoding $[0,1]$
4 Hour of the day, sinus encoding $[0,1]$
@@ 1 Hourly energy consumption [kWh] $\mathbb{R}_{+}$
We set cyclical features with a $\cos/\sin$ encoding similarly to [53] and introduce a binary flag to indicate weekends and public holidays. The dataset is created from proprietary data obtained through our collaboration with Hydro-Québec.
To follow the literature, we compare the WaDiRo-SCNN to a Gaussian process (GP), a WaDiRo linear regression, and a linear support vector regression (SVR). We refer readers to Appendix C for details on each model’s hyperparameter search space and the definition of the GP kernel. We limit the hyperparameter tuning to 15 runs per model to represent industrial computational constraints and use the tree-structured parzen estimator (TPE) algorithm proposed by [54] to navigate the search space. We have observed experimentally that TPE converges faster than simulated annealing to good solutions, which suits the limit on runs. The mean average error is used to guide the hyperparameter tuning. We acknowledge that these models are applied naively and that more sophisticated methods could be preferable for this specific real-world application: our goal is to highlight how the WaDiRo-SCNN compares to other models typically used in critical sectors on a real-world dataset and showcase its simplicity. We also test a simple physics-constrained WaDiRo-SCNN in which we enforce that energy consumption must be greater or equal to zero.
@@ 6.2.2 Numerical Results
Model WaDiRo-PCSCNN WaDiRo-SCNN GP Linear SVR WaDiRo linear reg.
@@
We observe that the WaDiRo-SCNN performs on par with the GP. Still, its performance is slightly above as it has the lower MAE between the two on 10 buildings out of 16 while having similar MSEs. We observe that the linear SVR has the worst performance of all models as it has a consistently lower performance than the WaDiRo linear regression. We observe no significant performance gap between the unconstrained and physics-constrained WaDiRo-SCNNs. We also remark that the physics-constrained WaDiRo-SCNN has the lowest number of constraint violations throughout the experiment. Additional figures, viz., the absolute errors and test prediction plots, are presented in Appendix C for completeness.
@@
We remark that both the GP and our models have different strengths. Our model can include hard physical constraints easily, doesn’t require making assumptions on the type of data seen in training, and is conservative by design. GPs offer empirical uncertainty bounds and can be defined for multiple outputs.
7 Closing remark
@@
In this work, we reformulate the $\operatorname{ReLU}$-SCNN training problem into its order-1 Wasserstein distributionally robust counterpart by using the $\ell_{1}$-loss function and by conditioning on the set of sampling vectors required by the SCNN. We demonstrate that the training problem has inherent out-of-sample performance guarantees. We show that our settings allow the inclusion of convex physical constraints, which are then guaranteed to be respected during training. We showcase the conservatism of our model in two experiments: a synthetic experiment with controlled data corruption and the real-world prediction of non-residential buildings’ hourly energy consumption.
@@
This work does have some limitations that must be highlighted. The WaDiRo-SCNN training problem is only defined for one output neuron, it is not yet parallelizable and, as most convex optimization problems, it takes an increasing amount of time to solve when augmenting its dimensionality, e.g., the maximal numbers of neurons in the hidden layers, the size of the training set, and the number of features. As such, adapting its formulation for graphics processing unit (GPU) acceleration, similarly to work done by [19], would be interesting. We remark that the integration of physical constraints has not been covered much as it will be investigated further in future works.
@@ Acknowledgments
This work was possible thanks to funding from the Fonds de recherche du Québec – Nature et technologies (FRQNT), the Natural Sciences and Engineering Research Council of Canada (NSERC), Mitacs Accelerate, and Hilo by Hydro-Québec. Special thanks to Salma Naccache, Bertrand Scherrer, Steve Boursiquot, Ahmed Abdellatif, and Odile Noël from Hilo.
@@ A Proof of Proposition 1
Our objective is to rewrite (WDR-SCNNT) in the form of (TWDRO). This proof is derived from results of [36] and is adapted here for completeness. Consider the $\ell_{1}$-norm loss function, we have $h_{\bm{\beta}}(\hat{\mathbf{z}})=|\hat{\mathbf{z}}^{\top}\bm{\beta}|$ and
@@
Considering the two linear optimization problems, we can find the equivalent dual problems A and B with dual variables $\lambda_{A},\ \lambda_{B}$, respectively.
@@
Because the objective function of both dual problems is 0, to have finite optimal values for both primals, (Dual-A) and (Dual-B) need to have a non-empty feasible set.
This implies that these two constraints must be respected:
@@
From this we find that
@@
${\bm{\beta}}(1-\lambda_{B})$ $=\bm{\theta}\ \text{with}\ \lambda_{B}\geq 0$
${\bm{\beta}}(1+\lambda_{A})$ $=\bm{\theta}\ \text{with}\ \lambda_{A}\leq 0.$
@@
For the $\ell_{1}$-loss, when $\|{\bm{\beta}}\|_{*}=\|{\bm{\beta}}\|_{\infty}$, the problem now becomes:
@@
For the $\ell_{2}$-loss, $\|{\bm{\beta}}\|_{*}$ is equal to the $\ell_{2}$-norm [37]. By the same process and noting that $a\geq 0$, because a norm is non-negative, we obtain the second equivalent problem:
@@ B.1 Hyperparameters and setting
The hyperparameters of each model are presented in Table 4. Note that all models are formulated with bias weights by default and that different regularization methodologies are cast as hyperparameters in this experiment.
Table 4: Detailled information on the synthetic experiment
@@
Norm in the order-1 Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Norm in the order-1 Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Number of neurons in hidden layers
@@
We sample only 2000 data points for each benchmark function before splitting them between training, validation, and training. For the benchmark functions that can be defined in dimension higher than two, viz., Ackley and Keane, we impose a dimension of four.
WaDiRo-SCNNs and SCNNs are modelled and optimized in cvxpy [55]. We then extract the weights and encode them into pytorch’s neural network module [56]. Similarly, the WaDiRo linear regression and the regularized linear regression are optimized with cvxpy, and the deep FNN is fully implemented with pytorch.
@@ B.2 Additionnal figures
Figure 4 presents the absolute values of the figures presented in Section 6.1.2.
@@
Figure 4: Absolute error of each synthetic experiment
@@ C.1 Hyperparameters and setting
First, we define the kernel used for the Gaussian process regression. Following our inspection of the data, we construct the GP kernel by adding a radial basis function (RBF) kernel to an exponential sine kernel to model periodicity and a white noise kernel to capture the possible noise. Both the GP and the linear SVR were implemented with scikit-learn [57].
@@
Table 5: Models’ hyperparameters for the building experiment
@@
Norm in the order-1 Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
Norm in the order-1 Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
Gaussian process length_scale
@@
Periodicity of the exponential sine kernel
Lenght-scale of the exponential sine kernel

### v2

@@ 8.1.2 Numerical Results
We now present and analyze the results of each synthetic experiment. Figure 3 shows the normalized error metrics of each model on each function. The normalization is done within each column so that 1.00 and 0.00 represent, respectively, the highest and lowest error between all models for a specific metric and function, e.g., the MAE on McCormick’s function. We propose this representation because different functions have patterns of different complexity and it highlights how models relatively compare to each other when facing the same level of pattern complexity. Nonetheless, the absolute values are also presented in Appendix B as a complement.
@@
The objective of Experiment A is to evaluate the ability of each model at learning over a highly adversarial training set while having the possibility to refine, indirectly through the validation phase, on a slightly out-of-sample distribution. This experiment was designed to see how different models’ regularization techniques may prevent them from overfitting on training outliers. As we can see, the WaDiRo-SCNN, regularized SCNN, and deep FNN perform similarly except when trying to learn Ackley where the WaDiRo-SCNN has the lowest test error. In this experiment, the linear models perform poorly compared to other models. We hypothesize that the validation phase is uncorrupted enough to let nonlinear models adjust to the non-convex functions during hyperparameter optimization. As expected the non-regularized SCNN leads to performance lower than its regularization counterpart in this particular setting.
@@ 8.2.1 Setting
Montréal’s continental humid climate delivers harsh winters and hot summers. As such, a significant amount of energy is consumed for both electric heating and air conditioning. Moreover, date and time features have a strong influence on non-residential buildings’ consumption patterns as their activities may change within days and weeks, e.g., opening hours, occupancy, and seasonal activities.
In this experiment, we seek to predict the year-round hourly energy consumption in kWh of 16 distinct non-residential buildings located in three adjacent distribution substations in Montréal, Québec, Canada. As we do not have any information on building activity, this task is done using only hourly consumption history and historical meteorological features. We use a 12-week history to train each model to predict the following week. We have approximately 2.5 years of collected data on each building. We proceed with a 12-week sliding window and collect the predictions for each new week. We use the data of the first year and a half for hyperparameter tuning and the data of the last year for performance assessment. The features and label are listed in Table 2.
@@
1 Day of the week, cosine encoding $[0,1]$
2 Day of the week, sine encoding $[0,1]$
3 Hour of the day, cosine encoding $[0,1]$
4 Hour of the day, sine encoding $[0,1]$
@@ 1 Hourly energy consumption [kWh] $\mathbb{R}_{+}$
We set cyclical features with a $\cos/\sin$ encoding similarly to (Moon et al. 2018) and introduce a binary flag to indicate weekends and public holidays. The dataset is created from proprietary data obtained through our collaboration with Hydro-Québec.
To follow the literature, we compare the WaDiRo-SCNN to a Gaussian process (GP), a WaDiRo linear regression, and a linear support vector regression (SVR). Both the GP and the linear SVR were implemented with scikit-learn (Pedregosa et al. 2011) while the WaDiRo-SCNN and WaDiRo linear regression have been implemented similarly to the previous experiment. We refer readers to Appendix C for details on each model’s hyperparameter search space and the definition of the GP kernel. We limit the hyperparameter tuning to 15 runs per model to represent industrial computational constraints and use the tree-structured Parzen estimator (TPE) algorithm proposed by (Bergstra et al. 2011) to navigate the search space. We have observed experimentally that TPE converges faster than simulated annealing to good solutions, which suits the limit on runs. The mean average error is used to guide the hyperparameter tuning. We also test a simple physics-constrained WaDiRo-SCNN in which we enforce that energy consumption must be greater or equal to zero.
With this experiment, we first aim to showcase how our models perform on real-world VPP data compared to the forecasting models actually used in critical industrial VPP applications. We also seek to demonstrate how the inclusion of hard physical constraints in the SCNN training affects the solution. We acknowledge that these models are applied naively and that more sophisticated methods could be preferable for this specific application.
@@ 8.2.2 Numerical Results
Model WaDiRo-PCSCNN WaDiRo-SCNN GP SVR WaDiRo lin. reg.
@@
We observe that the WaDiRo-SCNN performs on par with the GP as both models seem adequate to model nonlinear patterns. Still, its performance is slightly above as it has the lower MAE between the two on 10 buildings out of 16 while having similar RMSEs. Note that trained GPs may only converge to local optima. We observe that the linear SVR has the worst performance of all models as it has a consistently lower performance than the WaDiRo linear regression. We observe no significant performance gap between the unconstrained and physics-constrained WaDiRo-SCNNs. We also remark, as shown in Table 3, that the WaDiRo-PCSCNN has the lowest number of constraint violations throughout the experiment and that the WaDiRo-SCNN has the second lowest amount with yet 59.12% more. Additional figures, viz., the absolute errors and test prediction plots, are presented in Appendix C for completeness.
@@
We remark that both the GP and WaDiRo-SCNNs have different strengths. Our model can include hard physical constraints easily, does not require making assumptions on the type of data seen in training, and is conservative by design. GPs offer empirical uncertainty bounds and can be defined for multiple outputs.
8.3 Stability certification
We now design a final experiment to evaluate the stability of SCNNs trained with varying numbers of neurons and degrees of regularization.
8.3.1 Setting
We propose a grid search over the maximal number of neurons and the regularization parameter. We proceed on different popular real-world datasets from the UCI ML repository, namely the energy efficiency dataset from Tsanas and Xifara (2012) and the concrete compressive strength dataset from Yeh (1998). Additional tests on other real-world datasets are found on our GitHub page. The random seed for the sampling of $\mathcal{S}$ is fixed throughout the experiments, meaning that any instability that may arise from the sampling will be highlighted similarly for each model. We collect predictions in the SCNN form without mapping back to the non-convex SNN. We provide testing MAEs as well. Both the stability bounds and the MAEs are in the scale of the original dataset, but a standard scaler is used during training. As such, the inverse transform from the normalized outputs to the original scale may have an additional effect on the size of $L_{\varepsilon}$.
8.3.2 Numerical results
Experiment results are presented in Figures 5 and 6. We observe some general tendencies across the two datasets and the different models. The minimum and maximum of each scale are identified in the top left corner of each subfigure. First, the lowest test errors are achieved for high neuron counts and low regularization. Instability seems inversely correlated with the size of the regularization parameter. Thus, there seems to be a tradeoff between precision and stability. We note that some neurons count are anomalously less stable than others, e.g., 6 neurons in Figure 5, we believe that this is related to the population of sampling vectors used as the same outlying patterns emerge similarly for all models.
The experiment on the energy dataset demonstrates that neural networks can be highly unstable with lower regularization. Indeed, with the added effect of the scaling, $L_{\varepsilon}$ reaches the millions. For a high number of neurons and an intermediate regularization parameter, we note that there seems to be a sweet spot where test errors are low and stability is high.
Overall, this experiment validates experimentally that our WaDiRo training methodology operates as a regularizer.
(a) MAE test: LASSO-SCNN
(b) $L_{\varepsilon}$: LASSO-SCNN
(c) MAE test: WaDiRo-SCNN $\left(W_{\|\cdot\|_{1},1}\right)$
(d) $L_{\varepsilon}$: WaDiRo-SCNN $\left(W_{\|\cdot\|_{1},1}\right)$
(e) MAE test: WaDiRo-SCNN $\left(W_{\|\cdot\|_{2},1}\right)$
(f) $L_{\varepsilon}$: WaDiRo-SCNN $\left(W_{\|\cdot\|_{2},1}\right)$
Figure 5: Experiments on concrete
(a) MAE test: LASSO-SCNN
(b) $L_{\varepsilon}$: LASSO-SCNN
(c) MAE test: WaDiRo-SCNN $W_{\|\cdot\|_{1},1}$
(d) $L_{\varepsilon}$: WaDiRo-SCNN $W_{\|\cdot\|_{1},1}$
(e) MAE test: WaDiRo-SCNN $W_{\|\cdot\|_{2},1}$
(f) $L_{\varepsilon}$: WaDiRo-SCNN $W_{\|\cdot\|_{2},1}$
Figure 6: Experiments on energy
9 Closing remarks
@@
In this work, we reformulate the $\operatorname{ReLU}$-SCNN training problem into its tractable convex order-1 Wasserstein distributionally robust counterpart by using the $\ell_{1}$-loss function and by conditioning on the set of sampling vectors required by the SCNN. We demonstrate that the training problem has inherent out-of-sample performance guarantees. We show that our settings allow the inclusion of convex physical constraints, which are then guaranteed to be respected during training, and easy integration in mixed-integer convex post-training verification programs. These results showcase the strength and potential of convex learning for critical applications. We showcase the conservatism of our model in three experiments: a synthetic experiment with controlled data corruption, the real-world prediction of non-residential buildings’ hourly energy consumption, and two datasets from the UCI ML repository on which we certify model stability.
@@
This work does have some limitations that must be highlighted. The WaDiRo-SCNN training problem is only defined for one output neuron, it is not yet GPU parallelizable, and, as most convex optimization problems, it takes an increasing amount of time to solve when augmenting its dimensionality, e.g., the maximal numbers of neurons in the hidden layers, the size of the training set, and the number of features. As such, adapting its formulation for GPU acceleration or adapting the algorithm to a more efficient solving method , similarly to work done by Mishkin et al. (2022), Bai et al. (2023b) and more recently by Miria Feng and Pilanci (2024), would be highly interesting. We remark that the integration of physical constraints has not been covered much as it will be investigated further in future works on nonlinear optimal control.
@@ Acknowledgments
This work was possible thanks to funding from the Fonds de recherche du Québec (FRQ), the Natural Sciences and Engineering Research Council of Canada (NSERC), Mitacs Accelerate, and Hilo by Hydro-Québec. Special thanks to Salma Naccache and Bertrand Scherrer for their active support and the enriching discussions, as well as Steve Boursiquot, Ahmed Abdellatif, and Odile Noël from Hilo for making this project possible.
@@ A Proof of Proposition 1
Our objective is to rewrite (WDR-SCNNT) in the form of (TWDRO). This proof is derived from results of Chen and Paschalidis (2020) and is adapted here for completeness. Consider the $\ell_{1}$-norm loss function. We have $h_{\bm{\beta}}(\hat{\mathbf{z}})=|\hat{\mathbf{z}}^{\top}\bm{\beta}|$ and
@@
Considering the two linear optimization problems, we can find the equivalent dual problems A and B with dual variables $\lambda_{A}$ and $\lambda_{B}$, respectively:
@@
Because the objective function of both dual problems is 0, to have finite optimal values for both primals, (Dual-A) and (Dual-B) need to have a non-empty feasible set. This implies that these two constraints must be respected:
@@
From this we find that,
@@
${\bm{\beta}}(1-\lambda_{B})$ $=\bm{\theta}\ \text{where}\ \lambda_{B}\geq 0$
${\bm{\beta}}(1+\lambda_{A})$ $=\bm{\theta}\ \text{where}\ \lambda_{A}\leq 0.$
@@
For the $\ell_{1}$-loss, $\|{\bm{\beta}}\|_{*}=\|{\bm{\beta}}\|_{\infty}$, and the problem becomes:
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
For the $\ell_{2}$-loss, $\|{\bm{\beta}}\|_{*}$ is equal to the $\ell_{2}$-norm (Boyd and Vandenberghe 2004). By the same process and noting that $a\geq 0$ because a norm is non-negative, we obtain the second equivalent problem:
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@ B.1 Hyperparameters and setting
The hyperparameters of each model are presented in Table 4. Note that all models are formulated with bias weights by default and that different regularization methodologies are set as hyperparameters in this experiment.
Table 4: Hyperparameters of the synthetic experiment
@@
Norm of the Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Norm of the Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Number of neurons per layer
@@
We sample 2000 data points for each benchmark function before splitting them between training, validation, and testing. For the benchmark functions that can be defined in dimension higher than two, Ackley and Keane, we impose a dimension of four.
@@ B.2 Additionnal figures
Figure 7 presents the absolute values of the errors presented in Section 8.1.2. Label values for each function on the feature domain are needed to better understand the figure. We proceed to identify the max and min sampled values. For McCormick, the minimal value is -1.91 and the maximal value is 65.32. For PGandP, the minimal value is -3.12 and the maximal value is 6.78. For Keane, the minimal value is -9.13 and the maximal value is 0. For Ackley, the minimal value is 0 and the maximal value is 22.29.
Compared to the normalized results, McCormick is extremely polarized: some models have low nominal test errors while others have bigger ones. While the WaDiRo-SCNN performs well in Experiments A and B, it sees a medium performance in Experiment C. This is hidden by the extremely poor performance of the deep FNN in Experiment C. Figure 2 shows that McCormick has linear tendencies in the center of its domain but steep increases on its edges that linear models fail to capture.
Keane, with its small codomain, has a small error range. We observe that the SCNN with no regularization is the only one to fail in Experiment A, while both standard SCNN formulations fail comparatively to the other models in Experiment C.
One interesting takeaway from Figure 7 is that WaDiRo-SCNN, at its worst performance, is on par with linear regression models. We do see the standard SCNN formulations have a poorer performance than the linear models on Keane and the deep FNN has consistently worse performances in the corrupted setting from Experiment C.
@@
Figure 7: Absolute errors of each synthetic experiment
B.3 Computational time
The tests presented in Figure 8 were conducted on a computer with 64GB of RAM DDR5 and an AMD Ryzen 9 7950x3d CPU. The training was performed with the open-source solver Clarabel and was formulated in Python with cvxpy. As we can see, more neurons lead to smaller training errors while a greater dataset size increases the training error. Indeed, sampling more data points increases the difficulty of fitting them all correctly. As for the training time, to keep the computational time low, there is a tradeoff between the maximal number of neurons and the size of the dataset. This is expected partly because we do not yet have a GPU acceleration method for this training procedure. We also remark that a more direct implementation, e.g., by using a lower-level programming language and avoiding an interface like cvxpy for the constraint implementation, could also diminish the training time. As such, comparing these running times with methods implemented in C++ would be unfair.
(a) Training MAE
(b) Training time
Figure 8: Training error and time on Ackley
@@ C.1 Hyperparameters and setting
First, we define the kernel used for the Gaussian process regression. Following our inspection of the data, we construct the GP kernel by adding a radial basis function (RBF) kernel to an exponential sine kernel to model periodicity and a white noise kernel to capture the possible noise.
@@
Table 5: Hyperparameters of the building experiment
@@
orm used in Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
orm used in Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
GP length_scale
@@
Periodicity of the exp.-sine kernel
Lenght-scale of the exp.-sine kernel